OpenStudy (anonymous):

A bullet is fired horizontally with speed v from a rifle. For a short time t after leaving the rifle, the only force affecting its motion is gravity. The acceleration of free fall is g. Which expression gives the value of the horizontal distance travelled in time t/the vertical distance travelled in time t?

OpenStudy (anonymous):

a) vt/g b) v/gt c) 2vt/g d) 2v/gt

OpenStudy (anonymous):



OpenStudy (anonymous):

g affects only vertical motion not horizontal motion

OpenStudy (anonymous):

So v=gt?

OpenStudy (anonymous):

horiz distance = vt vert distance = v0t+(1/2)gt^2 = g*t^2 / 2 the ratio = \[\frac{vt}{gt^2\over2}=\frac{2v}{gt}\]

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

Okay. Thank you.